I made a mistake in the last entry. When I measured the frequency difference, I did not notice that two beams path were shifted at different orders from AOM.
One was +160MHz, another was -160 MHz. So I re measured the frequency difference again, with both beams shifted by 160MHz, the frequency difference between
the two cavities is 127 MHz. Without thermal adjust on the cavities,This falls into a hard-to-work-with frequency region with the current VCOs we have.
To lock both cavities at the same time with df = 127 MHz, for now, we use 0th order beam on RCAV's AOM, so that the beam's frequency is not shifted, and use a VCO [model's number] to drive ACAV's AOM at 63.6 MHz, so the frequency different between the two paths is 127 MHz, from this both cavities should be able to lock simultaeneously.
We should be able to see beat signal by this Tuesday.